OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 2x + y = 5 x + 3y = 0 –1 1 3 –3

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

you're welcome

OpenStudy (stamp):

2x + y = 5 x + 3y = 0 2x + y = 5 2(x + 3y = 0) 2x + y = 5 2x + 6y = 0 Subtract the two equations -5y = 5 Solve y = -1